OpenStudy (anonymous):

10p-2(3p-6)=4(5p-6)-10p

OpenStudy (anonymous):

distribute. Simplify. Solve.

OpenStudy (anonymous):

10p - 6p + 12 = 20p -24 -10p. 4p + 12 = 10p -24 36 = 6p p = 6

OpenStudy (anonymous):

10p-6p+12=20p-24-10p 4p+12=10p-24 36=6p p=6 yay